# How to Move Your Dragon: Text-to-Motion Synthesis For Large-Vocabulary Objects

### v1

@@ 3 Background: Hierarchical Skeletal Rig
Dynamic Features describe motion via a sequence of joint translations ($\mathbf{p}\in\mathbb{R}^{F\times J\times 3}$) and rotations ($\mathbf{R}\in\mathbb{R}^{F\times J\times D}$), where $F$ is the number of frames and $D$ the dimension of the rotation. $\mathbf{p}[f,j]$ and $\mathbf{R}[f,j]$ define the local translation and rotation of joint $j$ at frame $f$ relative to its parent. In this work, we focus on a simplified setting that considers only the temporal evolution of joint rotations, $\mathbf{R}$, as dynamic features.
@@ 4.2 Rig Augmentation
Moreover, we introduce three rig augmentation techniques to simulate various skeletal rig configurations found in real-world motion data, as illustrated in Figure 2. These techniques—Bone Length Augmentation, Joint Number Augmentation, and Rest Pose Resetting—expand the diversity of skeletal configurations in the Truebones Zoo dataset while preserving the original motion dynamics.
@@ 5.3 Handling Arbitrary Rest Pose
In hierarchical skeletal rig system, the dynamic features of motions are determined not only by the skeletal topology but also by the specific rest pose of each object. To ensure accurate modeling of motion dynamics, our framework incorporates explicit encoding of rest pose information. Specifically, the rest pose offset $\mathbf{P}_{\text{rest}}(j)\in\mathbb{R}^{3}$ for joint $j$ is transformed into a positional embedding $\text{RestPE}(j)\in\mathbb{R}^{D^{\prime}}$ by a simple MLP module, capturing its relative position and geometry within the skeleton in 3D space. These embeddings are added to the corresponding joint tokens:
@@ 5.4 Two-Stage Learning Approach
6 Experiments
6.1 Dataset & Evaluation Metrics
Following our two-stage learning approach, we first evaluate the pose synthesis model, then assess the final motion synthesis model. To evaluate pose synthesis, we aggregate all motion data for each object category and extract their poses. We then apply clustering, generating 30 distinct pose clusters per object. Three clusters are randomly selected as the test pose set, while the remaining clusters are used for training. For motion synthesis evaluation, we randomly select one motion per object for the test set, with the remaining motions used for training.
We evaluate the models using five widely adopted metrics to comprehensively assess their performance: Fréchet Inception Distance (FID) (Lee et al., 2019), R-Precision (Guo et al., 2022), Alignment (Guo et al., 2022), Coverage (Li et al., 2022), and Multimodality (Lee et al., 2019). FID measures the similarity between the feature distributions of generated and real data, with lower scores indicating more realistic outputs. R-Precision and Alignment evaluate how well generated samples match input conditions. R-Precision quantifies the proportion of correct matches, while Alignment measures cosine similarity, with higher values indicating better correspondence. Coverage assesses how much of the reference data is represented by the generated samples, considering a reference covered if its cosine similarity with generated data exceeds a threshold. Multimodality captures the diversity of generated outputs for a single input, with higher scores reflecting greater variability.
For evaluation, we train and utilize a pose-level encoder to compute all metrics instead of a motion-level encoder due to the limited availability of motion data. To ensure the pose encoder captures nuanced and meaningful representations of poses, we align its embedding space with that of the pretrained SigLIP image encoder, enabling the pose encoder to inherit the rich semantic knowledge learned during large-scale pretraining. We refer the reader to Table 1 (i.e. 1st row) and Figure 10 for the quantitative and qualitative evaluation of the learned pose encoder, respectively. Also find Appendix B.2 for additional details regarding the architecture and training protocol.
@@ 7.1 Generalization To Unseen Objects and Skeletons
We evaluate our model’s ability to generalize to novel objects with entirely different species and rig configurations. To this end, we obtain additional object meshes and armature data from external repositories such as SketchFab and use the curated text descriptions from Section 8 to generate motions. As shown in Figure 8, our model synthesizes coherent motions that align with reference dynamics, even for objects with previously unseen skeletal structures. This demonstrates its potential as a robust foundation for motion synthesis in open-vocabulary settings, expanding its applicability to diverse and novel object categories.
@@ 7.3 Generating Motions with Multi-Level Descriptions
8 Conclusion
We introduced a novel framework for text-driven motion synthesis across diverse object categories with heterogeneous skeletal structures. By augmenting the Truebones Zoo dataset with rich textual descriptions, introducing rig augmentation techniques, and extending motion diffusion models to dynamically adapt to arbitrary skeletal templates, our approach enables realistic, coherent motion generation for diverse objects. Experiments further demonstrate the potential of our method in generalization to unseen objects, long motion sequence generation, and controlled motion synthesis. We hope our work establishes a strong foundation for motion synthesis across diverse object categories, advancing many applications in 3D content creation.
9 Impact Statement

### v2

@@ 2.2 Non-Human Motion Synthesis
In contrast, CharacterMixer (Zhan et al., 2024) and SkinMixer (Nuvoli et al., 2022) propose rig-blending techniques that create hybrid characters and their motions constructed from multiple heterogeneous skeletons. While these methods demonstrate the potential to generalize motion across structurally diverse rigs, their reliance on motion retargeting limits flexibility in generating novel or diverse motions.
@@ 3 Background: Hierarchical Skeletal Rig
Dynamic Features describe motion via a sequence of joint translations ($\mathbf{p}\in\mathbb{R}^{F\times J\times 3}$) and rotations ($\mathbf{R}\in\mathbb{R}^{F\times J\times D}$), where $F$ is the number of frames and $D$ the dimension of the rotation. $\mathbf{p}[f,j]$ and $\mathbf{R}[f,j]$ define the local translation and rotation of joint $j$ at frame $f$ relative to its parent. In this work, we focus on a simplified setting that considers only the temporal evolution of rotations, $\mathbf{R}$, as dynamic features.
@@ 4.2 Rig Augmentation
Moreover, we introduce three rig augmentation techniques to simulate various skeletal rig configurations found in real-world motion data, as illustrated in Figure 2. These techniques, Bone Length Augmentation, Joint Number Augmentation, and Rest Pose Resetting, expand the diversity of skeletal configurations in the Truebones Zoo dataset while preserving the original motion dynamics.
@@ 4.3 High-Quality Multi-Level Text Annotation
For interested readers, detailed statistics on the dataset are provided in the appendix. These include the full list of animal categories in Truebones Zoo dataset (Table 3), word count histograms for each description level (Figure 10), and statistics on the most frequent verbs and nouns used in the annotations (Figure 11 and Figure 12, respectively).
@@ 5.3 Handling Arbitrary Rest Pose
In hierarchical skeletal rig system, the dynamic features of motions are determined not only by the skeletal topology but also by the specific rest pose of each object. To ensure accurate modeling of motion dynamics, our framework incorporates explicit encoding of rest pose information. Specifically, the rest pose offset $\mathbf{P}_{\text{rest}}(j)\in\mathbb{R}^{3}$ for joint $j$ is transformed into a positional embedding $\text{RestPE}(j)\in\mathbb{R}^{D^{\prime}}$ by a simple MLP module with sinusoidal encoding (Mildenhall et al., 2020), capturing its relative position and geometry within the skeleton in 3D space. These embeddings are added to the corresponding joint tokens:
@@ 5.4 Two-Stage Learning Approach
6 Experiments
All models were trained on a Linux system equipped with either an NVIDIA RTX A6000 (48GB) or A100 (40GB) GPU. The pose diffusion model required approximately 29GB of VRAM with a batch size of 512 over 400K iterations, completing training in roughly 30 hours. The motion diffusion model used about 38GB of VRAM with a batch size of 4 and sequence length of 90, trained for 1M iterations over approximately 4 days.
To ensure motion integrity of the augmented rigs, bone length adjustments are restricted to 0.8×0.8$\times$0.8 ×1.2×1.2$\times$1.2 × of the original and applied symmetrically when symmetry existed. Bone erasing was limited to distal appendages (toes, head tips, tail ends) and redundant spine/root bones. Under these constraints, we visually verified  10K augmented results and found them suitable for training.
6.1 Dataset & Evaluation Metrics
To evaluate the pose synthesis model, we aggregate all motion data for each object category and extract their poses. We then apply clustering, generating 30 distinct pose clusters per object. Three clusters are randomly selected as the test pose set, while the remaining clusters are used for training. For motion synthesis evaluation, we randomly select one motion per object for the test set, with the remaining motions used for training.
We evaluate the models using six automated metrics, along with a user study, to comprehensively assess their performance. The metrics include Fréchet Inception Distance (FID) (Lee et al., 2019), R-Precision (Guo et al., 2022), Alignment (Guo et al., 2022), Coverage (Li et al., 2022), Multimodality (Lee et al., 2019), and Motion Stability Index (MSI) (Kim et al., 2024). FID measures the similarity between the feature distributions of generated and real data, with lower scores indicating more realistic outputs. R-Precision and Alignment evaluate how well generated samples match input conditions. R-Precision quantifies the proportion of correct matches, while Alignment measures cosine similarity, with higher values indicating better correspondence. Coverage assesses how much of the reference data is represented by the generated samples, considering a reference covered if its cosine similarity with generated data exceeds a threshold. Multimodality captures the diversity of generated outputs for a single input, with higher scores reflecting greater variability. MSI measures temporal smoothness and stability, with higher scores indicating less jitter and stable motion over time. Finally, we conduct a user study to evaluate perceptual quality. Given rendered motion videos from each method, 30 participants compare the outputs in terms of realism, alignment with the input prompts, and overall preference. We aggregate the results across participants and report the average preference scores.
For automated evaluation, we train and utilize a pose-level encoder to compute all metrics instead of a motion-level encoder due to the limited availability of motion data. To ensure the pose encoder captures nuanced and meaningful representations of poses, we align its embedding space with that of the pretrained SigLIP image encoder, enabling the pose encoder to inherit the rich semantic knowledge learned during large-scale pretraining. We refer the reader to Table 1 (i.e. 1st row) and Figure 13 for the quantitative and qualitative evaluation of the learned pose encoder, respectively. Also find Appendix B.2 for additional details regarding the architecture and training protocol.
@@ 6.3 Text-to-Motion Synthesis
Table 2: User Preference and Smoothness (MSI) Scores.
Method Preference (%, ↑) MSI (×$\times 10^{3}$, ↑)
Ours 65.60 8.78
w/ GPT-Caption 2.27 7.30
w/o Rig Aug. 12.40 7.39
SO-MDMs 10.67 7.21
Retargeting 9.07 9.93
Finally, we report comparison results from a user study and a motion smoothness evaluation (measured by MSI), summarized in Table 2. Our method is consistently preferred by participants in terms of alignment with text prompts and overall motion quality, receiving the highest preference score by a substantial margin. Regarding smoothness, our approach achieves the highest score among all data-driven methods, indicating more temporally stable motions, while slightly trailing the retargeting method, which benefits from direct motion transfer but lacks flexibility to handle novel prompts. The results further highlight the effectiveness of our method in producing both perceptually preferable and plausible motions.
@@ 7.1 Generalization To Unseen Objects and Skeletons
We evaluate our model’s ability to generalize to novel objects with entirely different species and rig configurations.
To this end, we obtain additional object meshes and armature data from external repositories such as SketchFab and use the curated text descriptions from Section 8 to generate motions. As shown in Figure 8, our model synthesizes coherent motions that align with reference dynamics, even for objects with previously unseen skeletal structures. This demonstrates its potential as a robust foundation for motion synthesis in open-vocabulary settings, expanding its applicability to diverse and novel object categories.
@@ 7.3 Generating Motions with Multi-Level Descriptions
8 Discussion
We introduced a novel framework for text-driven motion synthesis across diverse object categories with heterogeneous skeletal structures. By augmenting the Truebones Zoo dataset with rich textual descriptions, introducing rig augmentation techniques, and extending motion diffusion models to dynamically adapt to arbitrary skeletal templates, our approach enables realistic, coherent motion generation for diverse objects. Experiments further demonstrate the potential of our method in generalizing to unseen objects, generating long motion sequences, and enabling controlled motion synthesis. We hope our work lays a strong foundation for motion synthesis across diverse object categories, advancing many applications in 3D content creation.
Despite its strengths, our method has a few limitations. First, for simplicity, our current implementation uses joint-wise Euler angle rotations only and ignores global root translation. However, this is not a fundamental limitation. Our model operates on motion tensors of shape $T\times J\times D$, where $D$ denotes joint features—in our case, rotations. This formulation can naturally be extended to include additional features such as global translation, relative translations for soft-constrained joints, or physically relevant signals like joint velocities or foot contact indicators. Including these signals may further improve temporal coherence, realism, and physical plausibility of synthesized motions.
Second, while our rig augmentation strategy enables generalization across diverse skeletal structures, the physical plausibility of augmented skeletons is limited. They may thus violate physical constraints (e.g. symmetry, balance, or anatomical consistency), limiting direct use in production-grade animation or simulation. Future work could improve plausibility via automated validation (e.g. foot-ground contact, joint velocity bounds, or end-effector stability), combined with rejection sampling or two-stage training. For example, pretraining on diverse augmented rigs for generalization, then fine-tuning on physically grounded data, may better balance flexibility and realism.
Third, while the method advances the field by offering a unified framework for motion synthesis across highly heterogeneous skeletal structures, it does not generalize well to human or open-vocabulary objects in zero-shot settings. This stems from the limited diversity, scale, and anatomical coverage of the training data: the Truebones Zoo dataset, though rich in non-human motions, lacks human-like skeletal structures and dynamics. Expanding to more balanced, large-scale motion corpora, such as Objaverse-XL (Deitke et al., 2024), could improve generalization to both human and novel object categories. This is a promising direction for future work, in line with recent advances in open-vocabulary dynamic mesh synthesis (Ren et al., 2023, 2024).
Acknowledgement
Wonkwang Lee was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2022-II220156, Fundamental research on continual meta-learning for quality enhancement of casual videos and their 3D metaverse transformation). Wonkwang Lee thanks Chaewon Kim, Hakyoung Lee, Howon Lee, Hyeongjin Nam, JungSeok Cho, Kangwook Lee, Moonwon Yu, Suekyeong Nam, and Seongjoon Yang for their insightful feedback during the early stages of this work. He also thanks Dongmin Park, Gihyun Kwon, Hyuck Lee, Hyunjin Kim, Hyunseung Kim, Joo Young Choi, Juhyeong Seon, Jinseo Jeong, Minkyu Kim, and Sehun Lee for their helpful discussions during the development of this work.
Impact Statement
@@ Acknowledgement
Appendix
This section presents additional details that are omitted from the main paper due to space constraints.
Figure 9: Rendered view images of a hamster’s pose from ten predefined camera viewpoints. We use these rendering images to annotate descriptions for each motion and to train the pose encoder.
@@ Appendix A Details on Data
For a visualization of word count distributions across description levels, see Figure 10. Statistics on the most frequent verbs and nouns are presented in Figure 11 and Figure 12, respectively.
@@
Table 3: The types of objects, the number of the 3D animations, and thier number of joints in Truebones Zoo dataset (Truebones, 2022).
Object Name # Animations Average # Joints
Alligator 21 21.00
Anaconda 18 27.00
Ant 16 34.00
Bat 11 43.00
Bear 30 62.00
Bird 19 53.00
Brown Bear 22 32.00
Buffalo 18 36.00
Buzzard 11 50.00
Camel 19 46.00
Cat 4 38.00
Centipede 9 64.00
Chicken 4 30.00
Komodo Dragon 9 50.00
Coyote 10 36.00
Crab 11 44.00
Cricket 11 43.00
Crocodile 12 34.00
Crow 10 23.00
Deer 21 37.00
Dog-1 42 46.00
Dog-2 36 46.00
Dragon 9 111.00
Eagle 12 39.00
Elephant 13 32.00
Fire Ant 25 35.00
Flamingo 4 34.00
Fox 13 35.00
Gazelle 18 35.00
Giant Bee 9 39.00
Goat 11 27.00
Hamster 5 38.00
Hermit Crab 12 55.00
Hippopotamus 10 36.00
Horse 29 63.00
Hound 12 40.00
Termite 9 52.00
Object Name # Animations Average # Joints
Jaguar 14 40.00
Great White Shark 10 16.00
King Cobra 11 17.18
Leopard 14 43.00
Lion 15 26.00
Lynx 14 33.00
Mammoth 12 36.00
Monkey 15 70.00
Ostrich 10 45.00
Parrot-1 14 56.00
Parrot-2 3 50.00
Pigeon 12 7.08
Piranha 11 22.00
Polar Bear 9 36.00
Baby Polar Bear 14 34.00
Pteranodon 13 34.00
Puppy 4 34.00
Reindeer 9 32.00
Velociraptor-1 10 33.00
Velociraptor-2 39 49.00
Velociraptor-3 14 50.00
Rat 7 13.00
Rhino 9 38.00
Roach 11 35.00
Sabre-Toothed Tiger 45 54.00
Sand Mouse 13 33.00
Scorpion-1 13 53.00
Scorpion-2 40 46.00
Skunk 8 29.00
Spider-1 33 57.00
Spider-2 18 51.00
Stegosaurus 10 35.00
Tyrannosaurus Rex 70 49.00
Triceratops 9 27.00
Toucan 9 15.00
Turtle 10 39.00
Tyrannosaurus 10 54.00
Figure 10: Word count histogram by description type.
Figure 11: Top 50 verbs and their counts in our annotated motion descriptions.
Figure 12: Top 50 nouns and their counts in our annotated motion descriptions.
@@ B.1 Generalized Motion Diffusion Models (G-MDMs)
Figure 13: Qualitative evaluation of the learned pose embedding space. Given a query pose of a sandmouse, we present rendered images of the top-1 retrieved poses from four distinct objects: coyote, horse, eagle, and crab. We display the file name of the motion file containing each retrieved pose. These results highlight the ability of the learned pose encoder to effectively embed similar poses from different animal objects into the same embedding space, capturing meaningful semantic relationships within the pose space.
@@ Appendix C Details on Coverage and Novelty Metrics
Figure 14: Quantitative comparisons of the models on the training set, evaluated using the coverage metric. While our model achieves comparable performance in fitting the training data compared to the version without rig augmentation, our full approach demonstrates significant gains in generalization (i.e. Figure 5).